OpenStudy (anonymous):

A soil supplement that weighs 19 lb contains iron, potassium, and mulch. The supplement contains sixteen times as much mulch as iron and twice as much potassium as iron. Find the amount of mulch in the soil supplement.

OpenStudy (jtug6):

1 sec

OpenStudy (anonymous):

k

OpenStudy (jtug6):

Alright so the first thing you'll want to do is write everything out thats given: It said 19 lb was equal to iron, potassium, and mulch, so we can write that as: 19 = i + p + m We also know that the supplement contains 16 times as much mulch as iron so we can write that as: 16m = i We also know theres twice as much potassium as there is iron so we write that as: 2p = i

OpenStudy (jtug6):

So now we need to solve for i for 16m = i and 2p = i once you find what value i is for those two plug them back into 19 = i + m + p then try from there

OpenStudy (anonymous):

so what should i do for step one

OpenStudy (jtug6):

Solve for i for 16m = i and 2p = i

OpenStudy (jtug6):

Because that gives you what m and p equal which you can then use back into 19 = i + m + p

OpenStudy (anonymous):

i+m+p=19

OpenStudy (jtug6):

yes

OpenStudy (jtug6):

but 16m = i and 2p = i so we need the values of m and p so we divide 16 for first and 2 of the other to get m and p

OpenStudy (anonymous):

8

OpenStudy (jtug6):

m = i/16 and p = i/2

OpenStudy (anonymous):

?

OpenStudy (jtug6):

so 19 = i + i/16 + i/2

OpenStudy (anonymous):

288/17

OpenStudy (jtug6):

Not quiet check again. i used 16 as common denom

OpenStudy (anonymous):

304/25

OpenStudy (jtug6):

yes!

OpenStudy (jtug6):

so now you know what iron is. it wants mulch right? so wasnt mulch 16m = i? and now we know i so....?

OpenStudy (anonymous):

ok now what

OpenStudy (jtug6):

16m = 12.16 so solve for m

OpenStudy (anonymous):

do i div

OpenStudy (jtug6):

yes. divide 16 across to solve for m

OpenStudy (anonymous):

1.3

OpenStudy (jtug6):

12.16/16 = ?

OpenStudy (anonymous):

0.76 sorry

OpenStudy (jtug6):

and there you have it :p

OpenStudy (anonymous):

that was wrong

OpenStudy (jtug6):

what were your choices then?

OpenStudy (anonymous):

No choice i have to find the Ib

OpenStudy (jtug6):

pretty sure it isnt wrong though. hmmm. did you include lb? does it want in lb or?

OpenStudy (anonymous):

No the system already has Ib for me

OpenStudy (anonymous):

just have to type the answer in

OpenStudy (jtug6):

odd. let me re-check thats definitley the way to solve i'll give you the answer once i re-do everything.

OpenStudy (anonymous):

ok

OpenStudy (jtug6):

I think i had my variables mixed up. Try 16. Youre not getting penalized for trying this right? I hope not T_T

OpenStudy (anonymous):

ok that was right thanks

OpenStudy (anonymous):

can u help me with one more question

OpenStudy (anonymous):

Find two consecutive even integers such that four times the first equals three times the second.

OpenStudy (jtug6):

I'm sorry. I shouldn't have made that issue but I read the question wrong. It was saying 16 times as much mulch as iron so instead of say 16m = i it was 16i = m. :X

OpenStudy (jtug6):

and same for potassium :p

OpenStudy (anonymous):

can u just tell the answer for this answer, so i can rate and give u points

OpenStudy (jtug6):

I'm not supposed to. I got in trouble last time so i'll just try to speed it up this time :p. 1 sec

OpenStudy (anonymous):

ok

OpenStudy (jtug6):

Ok so basically 4 times the first consecutive even integer is equal to 3 times the second consecutive even integer is basically saying: 4x = 3(x+2) So solve for x

OpenStudy (jtug6):

this will be really quick :p